AnimeGhoul8863:

The function H(t) = −16t2 + 75t + 25 shows the height H(t), in feet, of a baseball after t seconds. A second baseball moves in the air along a path represented by g(t) = 5 + 5.2t, where g(t) is the height, in feet, of the object from the ground at time t seconds. Part A: Create a table using integers 2 through 5 for the 2 functions. Between what 2 seconds is the solution to H(t) = g(t) located? How do you know? (6 points) Part B: Explain what the solution from Part A means in the context of the problem. (4 points)

AnimeGhoul8863:

@dude

dude:

 In order to solve for the y value, substitute in the x values (if you don't have a graph)

AnimeGhoul8863:

ok but how do i do this

AnimeGhoul8863:

how do i solve for x or y if neither of them are in the equation

AnimeGhoul8863:

@dude

dude:

x is 2, 3, 4, 5

AnimeGhoul8863:

ok

dude:

For example: \(H(2) = −16(2)^2 + 75(2) + 25\) and \(g(2) = 5 + 5.2(2)\)

dude:

This is only for when x=2

AnimeGhoul8863:

im still so confused

AnimeGhoul8863:

if x = 2,3,4,5 how do you know which one is going in the equation and which ones wont

AnimeGhoul8863:

i dont get this

dude:

It asks you in part A `Create a table using integers 2 through 5 for the 2 functions.` "Integers 2 through 5"

AnimeGhoul8863:

so what you gave is the answer for part A?

dude:

Part of it yes, we still have to solve for when x is 3, 4 and 5

AnimeGhoul8863:

but we didnt solve when it was 2 did we

AnimeGhoul8863:

Im sorry but im so lost

dude:

\(H(2) = −16(2)^2 + 75(2) + 25\\ -16\times 4+150+25\\ 64+150+25\\ H(2)=239\) and \(g(2) = 5 + 5.2(2)\\ 5+5.2\times 2\\ 5+10.4\\ g(2)=15.4\)

AnimeGhoul8863:

and do the same for 3 4 and 5

dude:

Right

AnimeGhoul8863:

ok ill try

AnimeGhoul8863:

h(3)=-16(3)^2+75(3)+25 h(3)=-16 x 3^2+75 x 3 +25 h(3)=-3^2 x 16+225+25 h(3)=250-3^2x16 3^2x16=144 h(3) 250-144 h(3) = 106 ^CORRECT?

dude:

That is right

AnimeGhoul8863:

h(4)=-16(4)^2+75(4)+25 h(4)=-16x4^2+75x4+25 h(4)=-4^2x16+300+25 h(4)=325-4^2x16 4^2x16=256 h(4)=325-256 H(4)=69

AnimeGhoul8863:

^CORRECT?

dude:

That is also correct

AnimeGhoul8863:

h(5)=-16(5)^2+75(5)+25 h(5)=-16x5^2+75x5+25 h(5)=-5^2x16+375+25 h(5)=400-5^2x 16 5x16=400 h(5)= 400-400= h(5)=0

AnimeGhoul8863:

^CORRECT?

AnimeGhoul8863:

Whats next

AnimeGhoul8863:

oh i forgot g

AnimeGhoul8863:

g(2) = 5 + 5.22 Add 10.22 Divide by both sides by 2 g(2)=5.11

AnimeGhoul8863:

^CORRECT

AnimeGhoul8863:

?

AnimeGhoul8863:

g(3) = 5 + 5.23 Add 10.23 Divide both sides by 3 g(3)=3.41 ^CORRECT?

AnimeGhoul8863:

g(4) = 5 + 5.24 Add 10.24 Divide both sides by 4 g(4)=2.56 ^CORRECT?

dude:

h(5) is correct, re check g(2)

AnimeGhoul8863:

g(5) = 5 + 5.25 Add 10.25 Divide both sides by 5 g(5)=2.05

dude:

You're supposed to multiply 5.2 and the x variable first for g(t) I showed an example for g(2)

dude:

There is no division required for g(t)

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude and \(g(2) = 5 + 5.2(2)\\ 5+5.2\times 2\\ 5+10.4\\ g(2)=15.4\) \(\color{#0cbb34}{\text{End of Quote}}\)

dude:

That is the example for g(2)

AnimeGhoul8863:

im confused now

AnimeGhoul8863:

i just looked it up https://www.symbolab.com/solver/algebra-calculator/g%5Cleft(2%5Cright)%3D5%2B5.22

dude:

That is assuming that g is a variable within the function, it is not important to the problem you only have to type whats after the g(#)

AnimeGhoul8863:

omg im sooooooo confused can you please explain more

AnimeGhoul8863:

i dont get thissssss

dude:

g(2) is just y when x equals 2 It doesn't mean anything that can help us in this problem You can completely ignore that when solving https://www.symbolab.com/solver/algebra-calculator/5%2B5.2%5Cleft(2%5Cright)

AnimeGhoul8863:

can you just help me find the answer for this i seriously am confused and i dont got how were going to answer this its 8PM im tired and i dont understand a word your saying

dude:

g(3) = 5 + 5.2*3 Multiply 5.2 and 3 \(\bf you~can~completely~ignore~the~g(3)for now\) 5+15.6 =20.6 That is g(3)

dude:

For g(4): g(4) = 5 + 5.2*4 Multiply 5.2 and 4 \(\bf you~can~completely~ignore~the~g(4)for~now\) 5+20.8 \(\bf =25.8\) This is g(4)

dude:

For g(5): g(5)=5+5.2*5 Multiply 5.2 and 5 \(\bf you~can~completely~ignore~the~g(5)for~now\) 5+26 \(\bf 31\) This is g(5)

AnimeGhoul8863:

so g(2)=15.4 g(3)=20.6 g(4)=25.8 g(5)=31 and h(2)=239 h(3)=106 h(4)=69 h(5)=0 Then what

AnimeGhoul8863:

Between what 2 seconds is the solution to H(t) = g(t) located? How do you know? how do we get 2 seconds to solution to = g(t)

dude:



AnimeGhoul8863:

ok

AnimeGhoul8863:

question is how do i put that in writing

dude:

https://www.desmos.com/calculator/6ugbhcc0y0 This is the graph it may help

dude:

The question asked for a table, so I gave you one

AnimeGhoul8863:

ok nvm ill figure it out ok so what does that make the answer

dude:

Between 4 and 5 seconds because we see that they both might have been at the same height and would've crossed each other

AnimeGhoul8863:

and that graph dosnt help it makes it confusing

AnimeGhoul8863:

ohhh ok so if thats the answer to part a lets work on part b

AnimeGhoul8863:

Explain what the solution from Part A means in the context of the problem.

dude:

So you see when x is 4, they haven't crossed each other? Now when you look at x=5 the line is already past the other equation They had intersected in between those points which was labeled `intersection!`

dude:

passed*

AnimeGhoul8863:

ok

dude:

So this represents that shortly after 4 seconds of being thrown both baseballs would have reach the same height

AnimeGhoul8863:

^answer for part B